Find the characteristics of the equation pq=xy
And determine the integral surface passes through the curve z=x , y=0
1
Expert's answer
2021-07-25T10:54:28-0400

We can also solve this question by method of separation of variables . We look for particular solution on the form -


"z(x,y)=f(x)g(y)"


"=(g\dfrac{df}{dx})(f\dfrac{dg}{dy})=xy\implies" "fdf=xdx" "\implies" "f={\pm}\sqrt{x^{2}+c_1}"

"gdg=ydy" "\implies" "g={\pm}\sqrt{y^{2}+c_2}"



So the function of z will become as -


"z(x,y)=" "{\pm}" "\sqrt{x^{2}+c_1}\sqrt{y^{2}+c_2}"


With condition z"(x,0)=x={\pm}\sqrt{(x^{2}+c_1)c_2}" "\implies" "c_1=0;c_2=1;" sign of "x"



"z(x,y)=x\sqrt{y^{2}+1}"


Solving with method of characteristics we get -


"=\dfrac{dx}{q}=\dfrac{dy}{p}=\dfrac{dz}{2pq}=\dfrac{dp}{y}=\dfrac{dq}{x}=\dfrac{dt}{1}"


"=x_0=s,y_0=0,z_0=s,p_0=1,q_0=0"


Solving the system leads to -


"x=Ae^{t}+Be^{-t}"

"y=Ce^{t}+De^{-t}"

"p=Ce^{2t}-De^{2t}"

"q=Ae^{2t}-Be^{2t}"

"z=ACe^{2t}-BDe^{2t}+E"


"=pq-xy=0" "\implies" "AB+BC=0"


with boundary conditions ; "p_0=1; q_0=0"


A"+B=s;\ C+D=0" ";AC-BD+E=s" ";C-D=1;A-B=0;AD+BC=0"


"A=B=E=\dfrac{s}{2}" and "C=-D=\dfrac{1}{2}\implies" "x=hcosht"

"y=sin(ht)"

"z=\dfrac{s}{2}(cosh(2t)+1)=scosh^{2}t"


"x^{2}-y^{2}s^{2}=s\implies" "s^{2}=\dfrac{x^{2}}{y^{2}+1}"


"z=\dfrac{x^{2}}{s}" "=x^{2}(\dfrac{+\sqrt{y^{2}+1}}{x})"


"z=x\sqrt{y^{2}+1}"